The accounting records for Portland Products report the following manufacturing costs for the past year.

Direct materials $ 340,000
Direct labor 261,000
Variable overhead 230,000

Production was 150,000 units. Fixed manufacturing overhead was $820,000.

For the coming year, costs are expected to increase as follows: direct materials costs by 20 percent, excluding any effect of volume changes; direct labor by 4 percent; and fixed manufacturing overhead by 10 percent. Variable manufacturing overhead per unit is expected to remain the same.


Required:

a. Prepare a cost estimate for a volume level of 120,000 units of product this year.
b. Determine the costs per unit for last year and for this year.

Cost Per Unit Last Year?

Cost Per Unit This Year?
Answer #1

Part A

Production unit 120000

Material cost per unit = 340000/150000 = 2.27

Direct labour cost per unit = 261000/150000 = 1.74

Variable overhead unit = 230000/150000 = 1.53

Fixed cost per unit = 820000/150000 = 5.47

Material cost increment 20% = 2.27 + 0.45 = 2.72

Labour cost increment 4% = 1.74 + 0.069 = 1.81

Fixed overhead increment 10% = 820000 + 82000 = 902000

Fixed overhead per unit = 902000/120000 = 7.52

Total cost per unit = 2.72 + 1.81 + 1.53 = 6.06

120000 * 6.06 = 727200

Total cost estimate = 727200 + 902000 = 1629200

Part B

Last year cost per unit = 2.27 + 1.74 + 1.53 + 5.47 = 11.01

This year cost per unit = 2.72 + 1.81 + 1.53 + 7.52 = 13.58